OpenStudy (anonymous):

Which is the factored form of x3 + 6x2 25x + 18 = 0? (Points : 4) (x + 1)(x 3)(x 6) = 0 (x + 1)(x 2)(x + 9) = 0 (x 1)(x 3)(x 6) = 0 (x 1)(x 2)(x + 9) = 0

OpenStudy (anonymous):

@andrewhaze So I took the test and am stuck on like 5 questions

OpenStudy (misty1212):

some stuff missing it looks like

OpenStudy (andrewhaze):

hold on

OpenStudy (anonymous):

(x - 1)(x -3)(x- 6) = 0 (x - 1)(x - 2)(x + 9) = 0 those are the last two

OpenStudy (andrewhaze):

. (x - 1)(x - 2)(x + 9) = 0

OpenStudy (anonymous):

Can you show me?

OpenStudy (andrewhaze):

@LeilaJudeh its (x - 1)(x - 2)(x + 9) = 0

OpenStudy (andrewhaze):

x*4+7x*2(x-1)-42x-x+42=x*4+7x*2(x-1)-42(x-1)+x=x*4+x+(x-1)(7x*2 -42)= =x(x*3 +1)+(x-1)(7x*2 -42)=x(x-1)(x*2+x-1)+(x-1)(7x*2-42)= =(x-1)(x(x*2+x-1)+7(x*2-6))

OpenStudy (anonymous):

okay